# No.2 Linggong Road, Dalian, Liaoning, 116024, P.R.China

### v1

@@ Abstract
Neural networks have emerged as a powerful paradigm for tasks in high energy physics, yet their opaque training process renders them as a black box. In contrast, the traditional cut flow method offers simplicity and interpretability but demands human effort to identify optimal boundaries. To merge the strengths of both approaches, we propose the Learnable Cut Flow (LCF), a neural network that transforms the traditional cut selection into a fully differentiable, data-driven process. LCF implements two cut strategies—parallel, where observable distributions are treated independently, and sequential, where prior cuts shape subsequent ones—to flexibly determine optimal boundaries. Building on this, we introduce the Learnable Importance, a metric that quantifies feature importance and adjusts their contributions to the loss accordingly, offering model-driven insights unlike ad-hoc metrics.  To ensure differentiability, a modified loss function replaces hard cuts with mask operations, preserving data shape throughout the training process. LCF is tested on six varied mock datasets and a realistic diboson vs. QCD dataset. Results demonstrate that LCF (1) accurately learns cut boundaries across typical feature distributions in both parallel and sequential strategies, (2) assigns higher importance to discriminative features with minimal overlap, (3) handles redundant or correlated features robustly, and (4) performs effectively in real-world scenarios. In diboson dataset, LCF initially underperforms boosted decision trees and multiplayer perceptrons when using all observables. However, pruning less critical features—guided by learned importance—boosts its performance to match or exceed these baselines. LCF bridges the gap between traditional cut flow method and modern black-box neural networks, delivering actionable insights into the training process and feature importance.
@@ 1 Introduction
In contrast to the cut flow method, machine learning is increasingly adopted across the field, proving effective in diverse applications: tagging jets with deep networks and jet charge 1511.05190 ; 1908.08256 ; 2012.08526 ; 2305.04372 , speeding up shower simulations with generative models 2210.06204 ; 2301.08128 ; 2305.17169 ; 2312.00042 ; 2401.02248 ; 2407.16704 ; 2410.21611 , enhancing unfolding for full-event reconstruction 2104.03036 ; 2105.09923 ; 2112.08180 ; 2203.16722 ; 2404.14332 , anomaly detection in complex LHC data 2305.15179 ; 2309.13111 ; 2310.06897 ; 2310.13057 ; 2311.17162 ; 2312.10009 , and suppressing pileup with attention-based methods 2107.02779 ; 2203.15823 .
@@
While machine learning excels in high-energy physics, its opacity has spurred efforts to align it with physical intuition and traditional methods. One approach embeds symmetry principles into neural networks, yielding architectures like Lorentz-equivariant autoencoders 2212.07347 , PELICAN networks 2211.00454 , energy flow networks 1810.05165 ; 2012.00964 , particle cloud jet tagging 1902.08570 , and others 2203.06153 ; 2212.00832 ; 2411.00446 . Another bridges neural networks with high-level observables, with ref. 1704.08249 leveraging ML to quantify jet information and saturate feature space, ref. 1710.01305 deriving novel jet observables, and ref. 1902.07180 automating observable construction. Additionally, researchers map low-level features to high-level observables 2010.11998 and craft interpretable anomaly detectors for jet substructure 2203.01343 , extracting discriminative insights in a human-readable form. For a comprehensive overview, see ref. 2102.02770 .
@@
To overcome the opacity of neural networks while retaining the clarity of traditional cut-based analysis, we introduce the Learnable Cut Flow (LCF). It is a neural network that fully emulates human cut-searching operations, ensuring its training process is entirely transparent and understandable. At bottom, we transform cuts into differentiable operations, such that training a neuron is equivalent to searching for an optimal cut. This match reflects LCF’s transparency and interpretability. For scenarios where signals lie in the middle or at both edges, we employ mask operations to split observable distributions into two parts around a specified center. With such learnable cuts, boundaries are automatically identified after training.
@@
Since observables vary in their separation power, contributing unevenly to classification, we introduce Learnable Importance. It is a set of softmax-normalized trainable weights that favors stronger observables while suppressing weaker ones. Unlike other ad-hoc metrics, it not only reflects the contribution fraction but also provides a robust way for selecting key observables to build the final cut flow.
To build the LCF model, we stack a series of layers. It starts with data preprocessing for standard normalization. Next comes the learnable importance layer. A split layer then separates data, passing each observable to its corresponding learnable cut. These cuts, matching the number of observables, process the data. Finally, a concatenation layer collects all cut results. During training, this structure learns the full cut flow. At inference, we select results based on importance—retaining cuts from high-importance observables and discarding the rest.
@@ 2.1 Learnable cuts
In high-energy physics, traditional cuts flows depend heavily on human effort. First, physicists select relevant observables. Next, each observable is visualized as a one-dimensional distribution. Then, boundaries are chosen by inspecting these distributions to separate signal from background events.
To formalize cuts, consider a dataset with $N$ event samples ($i$ indexes events) and $F$ observables ($j$ indexes observables). A traditional cut on $x_{ij}$ is:
$\hat{y}_{i}=\Theta\left(x_{ij}-c_{j}\right)$ (1)
$\hat{y}_{i}=1\rightarrow x_{ij}>c_{j}$ (2)
Here, $\Theta$ is the Heaviside step function, $c_{j}$ the cut’s boundary, and $\hat{y}_{i}$ the result. $\hat{y}_{i}=1$ means the event passes the cut $x_{ij}>c_{j}$ which is taken as a signal. We use the strictly greater than sign for clarity. This is a fundamental case where signal lies right of $c_{j}$ as a lower limit.
For the left case, with signal below the cut, the operation becomes:
$\hat{y}_{i}=\Theta\left(c_{j}-x_{ij}\right)$ (3)
$\hat{y}_{i}=1\rightarrow x_{ij}<c_{j}$ (4)
@@
To make cuts learnable and auto-updated, we replace the non-differentiable $\Theta$ with the logistic function $\sigma_{L}$, a smooth approximation common neural networks:
$\sigma_{L}(z)=\frac{1}{1+e^{-z}}$ (5)
@@
$\hat{y}_{i}=\sigma_{L}\left(x_{ij}-b_{j}\right)$ (6)
$\hat{y}_{i}>t\rightarrow x_{ij}>\sigma_{L}^{-1}(t)+b_{j}$ (7)
$\hat{y}_{i}=\sigma_{L}\left(b_{j}-x_{ij}\right)$ (8)
$\hat{y}_{i}>t\rightarrow x_{ij}<b_{j}-\sigma_{L}^{-1}(t)$ (9)
@@
$\hat{y}_{i}=\sigma_{L}\left(w_{j}x_{ij}-b_{j}\right)$ (10)
@@
$\hat{y}_{i}>t\rightarrow\begin{cases}x_{ij}>\text{boundary},&\text{if }\text{direction}>0\\ x_{ij}<\text{boundary},&\text{if }\text{direction}<0\end{cases}$ (11)
$\text{boundary}=\frac{1}{w_{j}}\left(\sigma_{L}^{-1}(t)+b_{j}\right)$ (12)
$\text{direction}=\text{sign}(w_{j})$ (13)
Here, the direction determines whether the signal lies above ($w_{j}>0$) or below ($w_{j}<0$) the boundary. This transforms a static cut to dynamic, learnable one, mimicking a neuron’s operation. Thus, training a neuron to tune $w_{j}$ and $b_{j}$ aligns with finding optimal boundary value—a clear connection from traditional cuts to neural networks.
@@
$\hat{y}_{ij}^{\text{lower}}$ $=\sigma_{L}\left(w_{j}^{\text{lower}}x_{ij}-b_{j}^{\text{lower}}\right)$ (14)
$\hat{y}_{ij}^{\text{upper}}$ $=\sigma_{L}\left(w_{j}^{\text{upper}}x_{ij}-b_{j}^{\text{upper}}\right)$
@@
$L_{ij}=\frac{1}{2}\left(L^{\text{lower}}_{ij}\operatorname{Mask}_{ij}^{\text{lower}}+L^{\text{upper}}_{ij}\operatorname{Mask}_{ij}^{\text{upper}}\right)$ (15)
$L_{ij}\left(y_{i},\hat{y}_{ij}\right)=-\left[y_{i}\log\left(\hat{y}_{ij}\right)+(1-y_{i})\log\left(1-\hat{y}_{ij}\right)\right]$ (16)
@@
$\text{Mask}^{\text{lower}}_{ij}=\begin{cases}1,&\text{if }x_{ij}<\text{center}_{j}\\ 0,&\text{otherwise}\end{cases}$ (17)
$\text{Mask}^{\text{upper}}_{ij}=\begin{cases}1,&\text{if }x_{ij}>\text{center}_{j}\\ 0,&\text{otherwise}\end{cases}$ (18)
To tie this together, we generalize all four cases (left, right, middle, and edge) using two sets of trainable parameters: $w_{j}^{\text{lower}}$, $b_{j}^{\text{lower}}$ for the lower range and $w_{j}^{\text{upper}}$, $b_{j}^{\text{upper}}$ for the upper range. These define a learnable cut for each observable, whether it’s a single-sided cut (left or right) or a double-sided one (middle or edge). For the simpler left and right cases, a single set of parameters can suffice, as shown in 11, but the two-set approach fully accommodates the more complex middle and edge scenarios. Based on later experimental results, we summarize the combination rules for the final learned cut:
@@ 2.2 Learnable importance
As the number of observables grows, determining their optimal combination becomes increasingly difficult. Unlike neural networks, which adapt to input features flexibly, traditional cut flows are sensitive to the order and selection of observables. Usually, highly-correlated observables are drop since they function similarly with each other. Those that have large overlap area between signal and background are also dropped since they don’t provide distinct characteristics between the two classes. We aim to let the LCF model to learn such an importance during the training process by providing a proper tool.
@@
$s^{\prime}_{j}=\sigma_{S}(\mathbb{s})_{j}=\frac{e^{s_{j}}}{\sum_{k=1}^{F}e^{s_{k}}}$ (19)
Here $\mathbf{s}=\left[s_{1},s_{2},\ldots,s_{F}\right]$ is the vector of trainable parameters across all observables, and $\sigma_{S}$ denotes the softmax function. While softmax is not the only option—any differentiable function ensuring $\sum_{j=1}^{F}s_{j}^{\prime}=1$ could work—its prevalence in machine learning makes it a natural choice.
@@
$\hat{y}_{ij}=\sigma_{L}\left(w_{j}x_{ij}^{\prime}-b_{j}\right)$ (21)
@@
$L_{ij}=-[y_{i}\log(\hat{y}_{ij})+(1-y_{i})\log(1-\hat{y}_{ij})]$ (23)
$\hat{y}_{ij}=\sigma_{L}(z_{ij})$ (24)
$z_{ij}=w_{j}x^{\prime}_{ij}+b_{j}$ (25)
$x^{\prime}_{ij}=s^{\prime}_{j}x_{ij}$ (26)
$s^{\prime}_{j}=\sigma_{S}(\boldsymbol{s})_{j}$ (27)
The gradient of $L$ with respect to $s_{k}$ reveal how importance scores adjust during training. We compute:
$\frac{\partial L}{\partial s_{k}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{F}\frac{\partial L_{ij}}{\partial\hat{y}_{ij}}\cdot\frac{\partial\hat{y}_{ij}}{\partial z_{ij}}\cdot\frac{\partial z_{ij}}{\partial s_{j}^{\prime}}\cdot\frac{\partial s_{j}^{\prime}}{\partial s_{k}}$ (28)
@@
$\frac{\partial L_{ij}}{\partial\hat{y}_{ij}}=\frac{\hat{y}_{ij}-y_{i}}{\hat{y}_{ij}\left(1-\hat{y}_{ij}\right)}$ (29)
$\frac{\partial\hat{y}_{ij}}{\partial z_{ij}}=\hat{y}_{ij}\left(1-\hat{y}_{ij}\right)$ (30)
$\frac{\partial z_{ij}}{\partial s_{j}^{\prime}}=w_{j}x_{ij}$ (31)
@@
$\frac{\partial s_{j}^{\prime}}{\partial s_{k}}=\begin{cases}s_{j}^{\prime}\left(1-s_{j}^{\prime}\right)&k=j\\ -s_{k}^{\prime}s_{j}^{\prime}&k\neq j\end{cases}$ (32)
@@
$\delta_{j}=\frac{1}{N}\sum_{i=1}^{N}\left(\hat{y}_{ij}-y_{i}\right)w_{j}x_{ij}$ (33)
@@
$\frac{\partial L}{\partial s_{k}}=\sum_{j=1}^{F}\delta_{j}\cdot\frac{\partial s_{j}^{\prime}}{\partial s_{k}}=s_{k}^{\prime}\left(\delta_{k}-\sum_{j=1}^{F}s_{j}^{\prime}\delta_{j}\right)$ (34)
This gradient drives the learning dynamics. Here, $\delta_{k}$ reflects the $k$-th observable’s misalignment between predictions and labels, while $\sum_{j=1}^{F}s_{j}^{\prime}\delta_{j}$ is a weighted baseline across all observables. If $\delta_{k}>$ baseline, the gradient is positive, reducing $s_{k}$ and thus $s_{k}^{\prime}$. This shrink the value of the observable towards zero and the $s_{k}^{\prime}$ outside the parenthesis reduce its gradient making it rely more on other influential ones. If $\delta_{k}<$ baseline, the gradient is negative, increasing $s_{k}^{\prime}$, boosting a reliable observable to keep the full range of it. This process embeds importance directly into training. Such a learned importance offers insights into each observable’s role and plays as a potential threshold for dropping negligible ones.

### v3

@@ Abstract
Neural networks have emerged as a powerful paradigm for tasks in high energy physics, yet their opaque training process renders them as a black box. In contrast, the traditional cut flow method offers simplicity and interpretability but requires extensive manual tuning to identify optimal cut boundaries. To merge the strengths of both approaches, we propose the Learnable Cut Flow (LCF), a neural network that transforms the traditional cut selection into a fully differentiable, data-driven process. LCF implements two cut strategies—parallel, where observable distributions are treated independently, and sequential, where prior cuts shape subsequent ones—to flexibly determine optimal boundaries. Building on this strategy, we introduce the Learnable Importance, a metric that quantifies feature importance and adjusts their contributions to the loss accordingly, offering model-driven insights unlike ad-hoc metrics. To ensure differentiability, a modified loss function replaces hard cuts with mask operations, preserving data shape throughout the training process. LCF is tested on six varied mock datasets and a realistic diboson vs. QCD dataset. Results demonstrate that LCF (1) accurately learns cut boundaries across typical feature distributions in both parallel and sequential strategies, (2) assigns higher importance to discriminative features with minimal overlap, (3) handles redundant or correlated features robustly, and (4) performs effectively in real-world scenarios. In the diboson dataset, LCF initially underperforms boosted decision trees and multilayer perceptrons when using all observables. LCF bridges the gap between traditional cut flow method and modern black-box neural networks, delivering actionable insights into the training process and feature importance. Source code and experimental data are available at https://github.com/Star9daisy/learnable-cut-flow.
arxiv: 2503.22498
@@ 1 Introduction
In contrast to the cut flow method, machine learning is increasingly adopted across the field, proving effective in diverse applications: tagging jets with deep networks and jet charge 1511.05190 ; 1908.08256 ; 2012.08526 ; 2305.04372 , speeding up shower simulations with generative models 2210.06204 ; 2301.08128 ; 2305.17169 ; 2312.00042 ; 2401.02248 ; 2407.16704 ; 2410.21611 , enhancing unfolding for full-event reconstruction 2104.03036 ; 2105.09923 ; 2112.08180 ; 2203.16722 ; 2404.14332 , anomaly detection in complex LHC data 2305.15179 ; 2309.13111 ; 2310.06897 ; 2310.13057 ; 2311.17162 ; 2312.10009 ; 2412.03747 , and suppressing pileup with attention-based methods 2107.02779 ; 2203.15823 .
@@
While machine learning excels in high-energy physics, its opacity makes the reinterpretation of the results far more involved, if not impossible in some cases, driving efforts to align such techniques with physical intuition and traditional methods. One approach embeds symmetry principles into neural networks, yielding architectures like Lorentz-equivariant autoencoders 2212.07347 , PELICAN networks 2211.00454 , energy flow networks 1810.05165 ; 2012.00964 , particle cloud jet tagging 1902.08570 , and others 2203.06153 ; 2212.00832 ; 2411.00446 . Another approach bridges neural networks with high-level observables, with ref. 1704.08249 leveraging ML to quantify jet information and saturate feature space, ref. 1710.01305 deriving novel jet observables, and ref. 1902.07180 automating observable construction. Additionally, researchers map low-level features to high-level observables 2010.11998 and craft interpretable anomaly detectors for jet substructure 2203.01343 , extracting discriminative insights in a human-readable form. For a comprehensive overview, see ref. 2102.02770 .
To overcome the opacity of neural networks while retaining the clarity of traditional cut-based analyses, we introduce the Learnable Cut Flow (LCF). It is a neural network that fully emulates human cut-searching operations, ensuring what it learns is entirely transparent and understandable. Recent efforts have also explored making cuts differentiable for optimization. In ref. Hance:2025haf , cuts are represented as bias parameters in a neural network, with losses that are customized for targeted efficiencies and smoothness. While both approaches share similarities in neural network structure, we adopt a different approach for deriving cut expression and calculating losses, emphasizing parallel and sequential strategies via mask operations, as well as a learnable importance mechanism for feature prioritization.
@@
At bottom, we transform cuts into differentiable operations, such that training a neuron is equivalent to searching for an optimal cut. This match reflects LCF’s transparency and interpretability. For scenarios where signals lie in the middle or at both edges of an observable distribution, we employ mask operations to split observable distribution into two parts around a specified center. With such learnable cuts, boundaries are automatically identified after training.
@@
Since observables vary in their separation power, contributing unevenly to classification, we introduce Learnable Importance. It is a set of softmax-normalized trainable weights that favors stronger observables while suppressing weaker ones. Unlike other ad-hoc metrics, it not only reflects the contribution fraction but also provides a robust way for selecting key observables to build the final cut flow. At inference, we select results based on importance—retaining cuts from high-importance observables and discarding the rest.
@@ 2.1 Learnable cuts
In high-energy physics, the traditional cut flow involves a series of manual steps. First, physicists select relevant observables. Next, each observable is visualized as a one-dimensional distribution. Then, boundaries are chosen by inspecting these distributions to separate signal from background events.
To formalize cuts, consider a dataset with $N$ events ($i$ indexes events) and $F$ observables ($j$ indexes observables). A traditional cut on $x_{ij}$ is
$\hat{y}_{i}=\Theta\left(x_{ij}-c_{j}\right),$ (1)
$\hat{y}_{i}=1\rightarrow x_{ij}>c_{j}.$ (2)
Here, $\Theta$ is the Heaviside step function, $c_{j}$ the cut’s boundary, and $\hat{y}_{i}$ the result. $\hat{y}_{i}=1$ means the event passes the cut $x_{ij}>c_{j}$, which is taken as a signal. We use the strictly greater than sign for clarity. This is a fundamental case where signal lies right of $c_{j}$ as a lower limit.
For the left case, with signal below the cut, the operation becomes
$\hat{y}_{i}=\Theta\left(c_{j}-x_{ij}\right),$ (3)
$\hat{y}_{i}=1\rightarrow x_{ij}<c_{j}.$ (4)
@@
To make cuts learnable and auto-updated, we replace the non-differentiable $\Theta$ with the logistic function $\sigma_{L}$, a smooth approximation commonly used with neural networks:
$\sigma_{L}(z)=\frac{1}{1+e^{-z}}.$ (5)
@@
$\hat{y}_{i}=\sigma_{L}\left(x_{ij}-b_{j}\right),$ (6)
$\hat{y}_{i}>t\rightarrow x_{ij}>\sigma_{L}^{-1}(t)+b_{j},$ (7)
$\hat{y}_{i}=\sigma_{L}\left(b_{j}-x_{ij}\right),$ (8)
$\hat{y}_{i}>t\rightarrow x_{ij}<b_{j}-\sigma_{L}^{-1}(t).$ (9)
@@
$\hat{y}_{i}=\sigma_{L}\left(w_{j}x_{ij}-b_{j}\right).$ (10)
@@
$\hat{y}_{i}>t\rightarrow\begin{cases}x_{ij}>\text{boundary},&\text{if }\text{direction}>0\\ x_{ij}<\text{boundary},&\text{if }\text{direction}<0\end{cases},$ (11)
$\text{boundary}=\frac{1}{w_{j}}\left(\sigma_{L}^{-1}(t)+b_{j}\right),$ (12)
$\text{direction}=\text{sign}(w_{j}).$ (13)
Here, the direction determines whether the signal lies above ($w_{j}>0$) or below ($w_{j}<0$) the boundary. This transforms a static cut to a dynamic, learnable one, mimicking a neuron’s operation. Thus, training a neuron to tune $w_{j}$ and $b_{j}$ aligns with finding the optimal boundary value—a clear connection from traditional cuts to neural networks.
For all results presented in Section 4, the threshold $t$ is set to 0.5, ensuring a balanced classification boundary consistent with standard binary classification practices. If we raise its value, through eq. (12) and eq. (13), $\sigma_{L}^{-1}(t)$ is increased since this function is monotonically increasing. Now the boundary change depends on the sign of $w_{j}$: if it’s positive (signal on the right), the boundary increases, shrinking the cut range; if it’s negative (signal on the left), the boundary decreases, similarly reducing the cut range. Therefore, higher $t$ values impose stricter cuts, potentially decreasing the number of events classified as signal, while lower $t$ values relax the cuts, expanding the signal region but risking increased background events.
@@
$\hat{y}_{ij}^{\text{lower}}$ $=\sigma_{L}\left(w_{j}^{\text{lower}}x_{ij}-b_{j}^{\text{lower}}\right),$ (14)
$\hat{y}_{ij}^{\text{upper}}$ $=\sigma_{L}\left(w_{j}^{\text{upper}}x_{ij}-b_{j}^{\text{upper}}\right).$
@@
$L_{ij}=\frac{1}{2}\left(L^{\text{lower}}_{ij}\operatorname{Mask}_{ij}^{\text{lower}}+L^{\text{upper}}_{ij}\operatorname{Mask}_{ij}^{\text{upper}}\right),$ (15)
$L_{ij}\left(y_{i},\hat{y}_{ij}\right)=-\left[y_{i}\log\left(\hat{y}_{ij}\right)+(1-y_{i})\log\left(1-\hat{y}_{ij}\right)\right],$ (16)
@@
$\text{Mask}^{\text{lower}}_{ij}=\begin{cases}1,&\text{if }x_{ij}<\text{center}_{j}\\ 0,&\text{otherwise}\end{cases},$ (17)
$\text{Mask}^{\text{upper}}_{ij}=\begin{cases}1,&\text{if }x_{ij}>\text{center}_{j}\\ 0,&\text{otherwise}\end{cases}.$ (18)
To tie this together, we generalize all four cases (left, right, middle, and edge) using two sets of trainable parameters: $w_{j}^{\text{lower}}$, $b_{j}^{\text{lower}}$ for the lower range and $w_{j}^{\text{upper}}$, $b_{j}^{\text{upper}}$ for the upper range. These define a learnable cut on each observable, whether it’s a single-sided cut (left or right) or a double-sided one (middle or edge). For the simpler left and right cases, a single set of parameters can suffice, as shown in eq. (11), but the two-set approach fully accommodates the more complex middle and edge scenarios. Based on later experimental results, we summarize the combination rules for the final learned cut:
@@ 2.2 Learnable importance
As the number of observables grows, determining their optimal combination becomes increasingly difficult. Unlike neural networks, which adapt to input features flexibly, traditional cut flows are sensitive to the order and selection of observables. Usually, highly-correlated observables are dropped since they function similarly with each other. Those that have a large overlap area between signal and background are also dropped since they don’t provide distinct characteristics between the two classes. We aim to let the LCF model to learn such an importance during the training process by providing a proper tool.
@@
$s^{\prime}_{j}=\sigma_{S}(\mathbf{s})_{j}=\frac{e^{s_{j}}}{\sum_{k=1}^{F}e^{s_{k}}}.$ (19)
Here $\mathbf{s}=\left[s_{1},s_{2},\ldots,s_{F}\right]$ is the vector of trainable parameters across all observables, $\sigma_{S}$ denotes the softmax function, and $F$ is the number of observables. While softmax is not the only option—any differentiable function ensuring $\sum_{j=1}^{F}s_{j}^{\prime}=1$ could work—its prevalence in machine learning makes it a natural choice.
@@
$\hat{y}_{ij}=\sigma_{L}\left(w_{j}x_{ij}^{\prime}-b_{j}\right).$ (21)
@@
$L_{ij}=-[y_{i}\log(\hat{y}_{ij})+(1-y_{i})\log(1-\hat{y}_{ij})],$ (23)
$\hat{y}_{ij}=\sigma_{L}(z_{ij}),$ (24)
$z_{ij}=w_{j}x^{\prime}_{ij}+b_{j},$ (25)
$x^{\prime}_{ij}=s^{\prime}_{j}x_{ij},$ (26)
$s^{\prime}_{j}=\sigma_{S}(\boldsymbol{s})_{j}.$ (27)
The gradient of $L$ with respect to $s_{k}$ reveals how importance scores adjust during training. We compute:
$\frac{\partial L}{\partial s_{k}}=\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{F}\frac{\partial L_{ij}}{\partial\hat{y}_{ij}}\cdot\frac{\partial\hat{y}_{ij}}{\partial z_{ij}}\cdot\frac{\partial z_{ij}}{\partial s_{j}^{\prime}}\cdot\frac{\partial s_{j}^{\prime}}{\partial s_{k}},$ (28)
@@
$\frac{\partial L_{ij}}{\partial\hat{y}_{ij}}=\frac{\hat{y}_{ij}-y_{i}}{\hat{y}_{ij}\left(1-\hat{y}_{ij}\right)},$ (29)
$\frac{\partial\hat{y}_{ij}}{\partial z_{ij}}=\hat{y}_{ij}\left(1-\hat{y}_{ij}\right),$ (30)
$\frac{\partial z_{ij}}{\partial s_{j}^{\prime}}=w_{j}x_{ij}.$ (31)
@@
$\frac{\partial s_{j}^{\prime}}{\partial s_{k}}=\begin{cases}s_{j}^{\prime}\left(1-s_{j}^{\prime}\right)&k=j\\ -s_{k}^{\prime}s_{j}^{\prime}&k\neq j\end{cases}.$ (32)
@@
$\delta_{j}=\frac{1}{N}\sum_{i=1}^{N}\left(\hat{y}_{ij}-y_{i}\right)w_{j}x_{ij},$ (33)
@@
$\frac{\partial L}{\partial s_{k}}=\sum_{j=1}^{F}\delta_{j}\cdot\frac{\partial s_{j}^{\prime}}{\partial s_{k}}=s_{k}^{\prime}\left(\delta_{k}-\sum_{j=1}^{F}s_{j}^{\prime}\delta_{j}\right).$ (34)
This gradient drives the learning dynamics. Here, $\delta_{k}$ reflects the $k$-th observable’s misalignment between predictions and labels, while $\sum_{j=1}^{F}s_{j}^{\prime}\delta_{j}$ is a weighted baseline across all observables. If $\delta_{k}>$ baseline, the gradient is positive, reducing $s_{k}$ and thus $s_{k}^{\prime}$. This shrinks the value of the observable towards zero and the $s_{k}^{\prime}$ outside the parenthesis reduces its gradient making it rely more on other influential ones. If $\delta_{k}<$ baseline, the gradient is negative, increasing $s_{k}^{\prime}$, boosting a reliable observable to keep the full range of it. This process embeds importance directly into training. Such a learned importance offers insights into each observable’s role and plays as a potential threshold for dropping negligible ones.